Could someone please tell me if these functions are odd, even or neither?
a. y = 2x^2
b. y = 2(x + 1)^2
c. y = 5x^1/4
d. y = x^2 - 4

I have a.even b. odd c. neither and d. even

thank you

Answer by lwsshak3(6463)

Could someone please tell me if these functions are odd, even or neither?
a. y = 2x^2
b. y = 2(x + 1)^2
c. y = 5x^1/4
d. y = x^2 - 4
**
if f(-x)=f(x), function is even
if f(-x)=-f(x), function is odd
..
a. y = 2x^2
f(-x)=2x^2=f(x), therefore, function is even
..
b. y = 2(x + 1)^2=2(x^2+2x+1)
f(-x)= 2(-x+1)^2=2(x^2-2x+1)
function neither odd or even
..
c. y = 5x^1/4
f(-x)=5(-x)^1/4
function neither odd or even, can't take an even root(4) of a negative number(-x)
..
d. y = x^2 - 4
f(-x)=(-x)^2-4=x^2-4=f(x), therefore, function is even